The number of numbers from 1-1000 which are neither divisible by 3 nor by 5 nor by 7 is

Grade:11

1 Answers

Saurabh Koranglekar
5 years ago
by 3= {1000/3= 333}
by 5= {1000/5 = 200}
by 7= {1000/7= 142}

by 3and 5 {1000/15=66}
by 3 and 7 {1000/ 21 =47 }
by 5 and 7 {1000/35 = 25 }

by 3 and 5 and 7 {1000/105 = 9 }

1000 – 333 – 200 – 142+66+47+25-9 = 454